Two blocks A and B of mass 1kg and 2 kg are connected together by means of a spring and are resting on a horizontal frictionless table. The blocks are then pulled apart so as to stretch the spring and then released. Find the ratio of their :


a) speed.                                           b) momentum                                            c) konetic energy at any instant.

Grade:11

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Arnab,

 

As the intial momentum of the system is zero (as they are relased from rest), and no external force acts on the system, the net momentum of the system should be 0 at any instant in time.

 

Let us say v1 and v2 are the speeds of the mass 1Kg and 2Kg respectively,

 

So 1*v1 = 2*v2

or v1/v2 = 2/1

And ratio of their momentum is 1.

Ratio of KE = 1*v1^2/2*v2^2 = 2.

 

Hope this helps.

 

Regards,

Ashwin (IIT Madras).